### v1

@@ 4.1 Off-policy Gradient
$M=1$ $M=5$ $M=10$
$\delta_{\textsc{mc}}$ $6.76\pm 0.16$ $6.76\pm 0.07$ $6.76\pm 0.05$
$\delta_{\textsc{ht}}$ $6.76\pm 0.16$ $6.76\pm 0.07$ $6.76\pm 0.05$
$\delta_{\textsc{cv}},{\scriptstyle\alpha=1}$ $6.28\pm 2.54$ $6.28\pm 1.13$ $6.28\pm 0.79$
$\delta_{\textsc{cv}}$ $6.76\pm 0.16$ $6.76\pm 0.07$ $6.76\pm 0.05$
@@
Therefore, the Monte Carlo estimator using samples $\boldsymbol{Y}^{(1)},\cdots,\boldsymbol{Y}^{(M)}\mathrel{\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}}p_{\boldsymbol{\theta}_{\text{old}}}$ is
@@
$\begin{split}\textsc{kl}(p_{\boldsymbol{\theta}}\mid\mid q)=\operatorname*{\mathbb{E}}_{\boldsymbol{Y}\sim p_{\boldsymbol{\theta}_{\text{old}}}}\Bigg{[}\frac{p_{\boldsymbol{\theta}}(\boldsymbol{Y})}{p_{\boldsymbol{\theta}_{\text{old}}}(\boldsymbol{Y})}\lim_{N\rightarrow\infty}\sum_{n=1}^{N}\\ \operatorname*{\mathbb{E}}_{\overline{Y}_{n}}\left[\log\frac{\vec{p}_{\boldsymbol{\theta}}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}_{<n})}{\vec{q}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}_{<n})}\right]\Bigg{]}.\end{split}$ (28)
Therefore, the Rao–Blackwellized Monte Carlo estimator using samples $\overline{\boldsymbol{Y}}^{(1)},\cdots,\overline{\boldsymbol{Y}}^{(M)}\mathrel{\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}}p_{\boldsymbol{\theta}_{\text{old}}}$ is
$\begin{split}\delta_{\textsc{rb}}^{\text{old}}=\frac{1}{M}\sum_{m=1}^{M}\frac{p_{\boldsymbol{\theta}}(\boldsymbol{Y}^{(m)})}{p_{\boldsymbol{\theta}_{\text{old}}}(\boldsymbol{Y}^{(m)})}\lim_{N\rightarrow\infty}\sum_{n=1}^{N}\quad\quad\\ \operatorname*{\mathbb{E}}_{\overline{Y}_{n}}\left[\log\frac{\vec{p}_{\boldsymbol{\theta}}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}_{<n}^{(m)})}{\vec{q}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}_{<n}^{(m)})}\right].\end{split}$ (29)
@@
Figure 2: Standard deviation of different KL estimators across various prompts in the IMDB evaluation set, displayed on a logarithmic scale. The KL divergence is measured between a GPT-2 model before and after direct preference optimization. The $\delta_{\textsc{cv}}$ estimator with $\alpha=1$ shows considerable instability for certain prompts, whereas the Rao–Blackwellized estimator consistently yields the lowest standard deviation, indicating more stable estimates.
@@ 5.1 Understanding the KL Estimators
In Fig. 1, we visualize the KL estimates for three different prompts: (left) a positive prompt, (middle) a neutral prompt, and (right) a negative adversarial prompt. The traces represent the average estimates from all the repetitions, while the shaded regions indicate the standard deviation. Except $\delta_{\textsc{cv}},\alpha=1$, all other estimators are unbiased and consistent. We observe that the $\delta_{\textsc{cv}}$ estimator exhibits a noticeable bias and high standard deviation when $\alpha=1$, i.e., when it is not set to its optimal value. The bias arises from numerical instability during the computation of $g(\boldsymbol{Y})=\frac{q(\boldsymbol{Y})}{p(\boldsymbol{Y})}-1$. The high variance is due to large values of $\mathop{\mathrm{Var}}\left[{\normalcolor g(\boldsymbol{Y})}\right]$. Specifically, for certain prompts, $\mathop{\mathrm{Var}}\left[{\normalcolor g(\boldsymbol{Y})}\right]$ can be unbounded. As the sample size increases, the chance of sampling from the tail of $p_{\boldsymbol{\theta}}$ also increases. These tail samples often correspond to negative movie reviews that had a high probability under the language model prior to fine-tuning, i.e., $q$, leading to extremely large values of $g(\boldsymbol{Y})$ and, consequently, a high standard deviation. This effect indeed depends on the prompt and is particularly pronounced for neutral and adversarial prompts.
@@
Figure 4: In comparison to models that are trained under the MC estimator, models that are trained with the RB estimator appear on the overall Pareto front $76\%$ of the time. We visualize the Pareto front for each estimator and show the fine-tuning runs that do not end up on the Pareto front with a lower opacity. RB estimator is more effective compared to the MC estimator in constraining the model to maintain a low KL while achieving reasonable rewards.
@@ Acknowledgements
We thank Ahmad Beirami for the insightful discussions throughout the course of this project. We also thank Alexander K. Lew for the valuable feedback on a draft of this paper. Afra Amini is supported by the ETH AI Center doctoral fellowship.
@@ Appendix A Horvitz–Thompson Estimation
Note that $\{\boldsymbol{Y}^{(m)}\}_{m=1}^{M}\mathrel{\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}}p$. Next, we compute the bias and variance of this estimator as follows:
@@ Appendix D Connection to Bregman Divergence
Bregman divergence (Bregman, 1967) is a measure of the difference between a convex function and its tangent plane. Let $F:\Omega\rightarrow\mathbb{R}$ be a strictly convex function, where $\Omega$ is a convex set. Bregman divergence associated with $F$ is defined as
@@
$D_{F}(p,q)$ =$\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\sum_{\boldsymbol{y}\in\Sigma^{*}}p(\boldsymbol{y})\log p(\boldsymbol{y})-q(\boldsymbol{y})\log q(\boldsymbol{y})-\big{(}\log q(\boldsymbol{y})+1\big{)}\big{(}p(\boldsymbol{y})-q(\boldsymbol{y})\big{)}$ (35a)
@@ F.1 Variance Reduction Proof
Let $\nabla\delta_{\textsc{mc}}^{N}$ be the length-conditioned MC estimator of the gradient defined in Eq. 43 and $\nabla\delta_{\textsc{rb}}^{N}$ the length-conditioned RB estimator of the gradient defined in Eq. 44. Define $G_{N}\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\operatorname*{\mathbb{E}}[\nabla\delta_{\textsc{mc}}^{N}]=\operatorname*{\mathbb{E}}[\nabla\delta_{\textsc{rb}}^{N}]$. We have
@@ F.2 A Note on Rao–Blackwellizing KL in Trust Region Algorithms
where $\boldsymbol{Y}^{(1)},\cdots,\boldsymbol{Y}^{(M)}\mathrel{\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}}p_{\boldsymbol{\theta}_{\text{old}}}$. Notably, the expected value of the above estimator is not equal to $\textsc{kl}(p_{\boldsymbol{\theta}}\mid\mid q)$ and is

### v2

@@ 4.1 Off-policy Gradient
Therefore, the Monte Carlo estimator using samples $\boldsymbol{Y}^{(1)},\cdots,\boldsymbol{Y}^{(M)}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny i.i.d.}}}{{\sim}}}p_{\boldsymbol{\theta}_{\text{old}}}$ is
@@
$\textsc{kl}(p_{\boldsymbol{\theta}}\mid\mid q)=\operatorname*{\mathbb{E}}_{\boldsymbol{Y}\sim p_{\boldsymbol{\theta}_{\text{old}}}}\Bigg{[}\frac{p_{\boldsymbol{\theta}}(\boldsymbol{Y})}{p_{\boldsymbol{\theta}_{\text{old}}}(\boldsymbol{Y})}\lim_{N\rightarrow\infty}\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{\overline{Y}_{n}}\left[\log\frac{\vec{p}_{\boldsymbol{\theta}}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}_{<n})}{\vec{q}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}_{<n})}\right]\Bigg{]}.$ (28)
Therefore, the Rao–Blackwellized Monte Carlo estimator using samples $\overline{\boldsymbol{Y}}^{(1)},\cdots,\overline{\boldsymbol{Y}}^{(M)}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny i.i.d.}}}{{\sim}}}p_{\boldsymbol{\theta}_{\text{old}}}$ is
$\delta_{\textsc{rb}}^{\text{old}}=\frac{1}{M}\sum_{m=1}^{M}\frac{p_{\boldsymbol{\theta}}(\boldsymbol{Y}^{(m)})}{p_{\boldsymbol{\theta}_{\text{old}}}(\boldsymbol{Y}^{(m)})}\lim_{N\rightarrow\infty}\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{\overline{Y}_{n}}\left[\log\frac{\vec{p}_{\boldsymbol{\theta}}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}_{<n}^{(m)})}{\vec{q}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}_{<n}^{(m)})}\right].$ (29)
@@ 5.1 Understanding the KL Estimators
$M=1$ $M=5$ $M=10$
$\delta_{\textsc{mc}}$ $6.76\pm 0.16$ $6.76\pm 0.07$ $6.76\pm 0.05$
$\delta_{\textsc{ht}}$ $6.76\pm 0.16$ $6.76\pm 0.07$ $6.76\pm 0.05$
$\delta_{\textsc{cv}}$ $6.28\pm 2.54$ $6.28\pm 1.13$ $6.28\pm 0.79$
$\delta_{\textsc{cv}1}$ $6.76\pm 0.16$ $6.76\pm 0.07$ $6.76\pm 0.05$
@@
In Fig. 1, we visualize the KL estimates for three different prompts: (left) a positive prompt, (middle) a neutral prompt, and (right) a negative adversarial prompt. The traces represent the average estimates from all the repetitions, while the shaded regions indicate the standard deviation. Except $\delta_{\textsc{cv}},\alpha=1$, all other estimators are unbiased and consistent. We observe that the $\delta_{\textsc{cv}}$ estimator exhibits a noticeable bias and high standard deviation when $\alpha=1$, i.e., when it is not set to its optimal value. The bias arises from numerical instability during the computation of $g(\boldsymbol{Y})=\frac{q(\boldsymbol{Y})}{p(\boldsymbol{Y})}-1$. The high variance is due to large values of $\mathop{\mathrm{Var}}\left[{\normalcolor g(\boldsymbol{Y})}\right]$. Specifically, for certain prompts, $\mathop{\mathrm{Var}}\left[{\normalcolor g(\boldsymbol{Y})}\right]$ can be unbounded.
Figure 3: Standard deviation of different KL estimators across various prompts in the IMDB evaluation set, displayed on a logarithmic scale. The $\delta_{\textsc{cv}}$ estimator with $\alpha=1$ shows considerable instability for certain prompts, whereas the $\delta_{\textsc{rb}}$ estimator consistently yields the lowest standard deviation, indicating more stable estimates.
As the sample size increases, the chance of sampling from the tail of $p_{\boldsymbol{\theta}}$ also increases. These tail samples often correspond to negative movie reviews that had a high probability under the language model prior to fine-tuning, i.e., $q$, leading to extremely large values of $g(\boldsymbol{Y})$ and, consequently, a high standard deviation. This effect indeed depends on the prompt and is particularly pronounced for neutral and adversarial prompts.
@@ 5.2 KL Estimation and RLHF Training Dynamics
Figure 4: Compared to MC-trained models, RB-trained models appear on the Pareto front $76\%$ of the time. We plot the reward vs. KL Pareto front for each estimator to illustrate how RB more effectively constrains KL while maintaining competitive rewards.
@@ Acknowledgements
We thank Ahmad Beirami and Cristina Pinneri for the insightful discussions throughout the course of this project. We also thank Alexander K. Lew for the valuable feedback on a draft of this paper. Afra Amini is supported by the ETH AI Center doctoral fellowship.
@@ Appendix A Horvitz–Thompson Estimation
Note that $\{\boldsymbol{Y}^{(m)}\}_{m=1}^{M}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny i.i.d.}}}{{\sim}}}p$. Next, we compute the bias and variance of this estimator as follows:
@@ Appendix D Connection to Bregman Divergence
Bregman divergence (Bregman, 1967) is a measure of the difference between a convex function and its tangent plane. Let $F\colon\Omega\rightarrow\mathbb{R}$ be a strictly convex function, where $\Omega$ is a convex set. Bregman divergence associated with $F$ is defined as
@@
$D_{F}(p,q)$ =$\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\sum_{\boldsymbol{y}\in\Sigma^{*}}p(\boldsymbol{y})\log p(\boldsymbol{y})-q(\boldsymbol{y})\log q(\boldsymbol{y})-\big{(}\log q(\boldsymbol{y})+1\big{)}\big{(}p(\boldsymbol{y})-q(\boldsymbol{y})\big{)}$ (35a)
@@ F.1 Variance Reduction Proof
Let $\nabla\delta_{\textsc{mc}}^{N}$ be the length-conditioned MC estimator of the gradient defined in Eq. 43 and $\nabla\delta_{\textsc{rb}}^{N}$ the length-conditioned RB estimator of the gradient defined in Eq. 44. Define $G_{N}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\operatorname*{\mathbb{E}}[\nabla\delta_{\textsc{mc}}^{N}]=\operatorname*{\mathbb{E}}[\nabla\delta_{\textsc{rb}}^{N}]$. We have
@@ F.2 A Note on Rao–Blackwellizing KL in Trust Region Algorithms
where $\boldsymbol{Y}^{(1)},\cdots,\boldsymbol{Y}^{(M)}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny i.i.d.}}}{{\sim}}}p_{\boldsymbol{\theta}_{\text{old}}}$. Notably, the expected value of the above estimator is not equal to $\textsc{kl}(p_{\boldsymbol{\theta}}\mid\mid q)$ and is